The following data were obtained: 525 nm

Concentration Absorbance
0.000008 0.005
0.000016 0.010
0.00004 0.040
0.00008 0.070

Plot the data and determine the trendline (slope and y-intercept)

Use Beer's law to calculate molar absorptivity for KMnO4 given the cell width to be 1 cm
Answer #1

From the graph we can see that at, y= 0.028 x= 0.000032 and at y=0.048 x=0.000054 and when x=0 y=0 as well.

So the y intercept is 0.

And the slope, dy/dx= (0.048-0.028)/(0.000054-0.000032) = 0.020/0.000022 = 909

According to Beer's law, A = εcl

Where ε is the molar absorptivity, c is the concentration of the sample, l is the path length in which light travels and A is the light absorbed.

From graph we can see that,

At c= 0.000054 the absorbance A= 0.048

Let, for a cell of c=1cm the Molar absorptivity ε=a.

So, 0.048=0.000054x1xa

or, a= 0.048/0.000054

or, a= 889

So the molar absorptivity of the sample is 889 mole-1Lcm-1